A transverse wave on a taut string is modeled with the wave function y(x, t) = Asin(kx-wt)=(0.2m)sin(6.28m-1 x – 1.57s-1 t) Find the: (a) amplitude, (b) wavelength, (c) period, (d) speed of the wave.
1
Expert's answer
2021-10-02T14:55:54-0400

The general equation for the transverse wave is given as follows:


"y(x,t) = A\sin \left(\dfrac{2\pi}{\lambda}x - 2\pi ft \right)"

where "A" is the amplitude, "\lambda" is the wavelength, "f" is the frequency of the wave, "v = \lambda f" is its speed and "T = 1\/f" is its period.

Comparing this expression with the given one, find:


"A = 0.2m\\
\dfrac{2\pi}{\lambda} = 6.28\Rightarrow\lambda \approx 1m\\
2\pi f = 1.57\Rightarrow f\approx 0.25Hz\\
T = 1/f \approx 4s\\
v = \lambda f = 0.25m/s"

Answer. a) 0.2m, b) 1m, c) 4s, d) 0.25 m/s.